A Complete Bibliography of Publications in the The Natural Philosopher

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

29 September 2023 Version 1.02

Title word cross-reference

111 [17]. 127 [16]. 1963 [14, 15].

2 [16].

Affairs [14, 15]. Ages [14, 15]. Annalen [18]. Annotated [4]. Aristotle [5]. Assistant [4].

Being [4]. betreffenden [13]. Bibliographic [11]. Blaisdell [17, 14, 15, 16]. Book [14, 15, 16].

chapter [5]. Charge [8]. Chemistry [3]. Company [16]. Concept [18]. Concepts [8]. Control [1]. Current [8].

Daniel [17, 16]. Devons [4]. Devoted [14, 15]. Duality [9]. edited [17, 16]. Einstein [18, 6, 9]. Eleatic [12]. Electric [8]. Electrical [7]. Electricity [11]. Entropy [2]. Erzeugung [13]. Faraday [11]. First [4, 5, 6]. foundation [5]. General [3]. Georges [10]. Georges-Louis [10]. German [13]. Gershenson [17, 16]. Gesichtspunkt [13]. Gravitational [10]. Great [7]. Greenberg [17, 16]. Here [4]. heuristic [13]. heuristischen [13]. History [14, 15]. Human [14, 15]. III [17]. Influence [14, 15]. Introduction [3]. Laboratory [4]. Lavoisier [3]. Liber [5]. Lichtes [13]. light [13]. Lord [4]. Louis [10]. Machine [7]. Marum [7]. Metaphysica [5]. Natural [17, 14, 15, 16]. Paper [18, 6]. Papers [11]. Part [8]. Particle [9]. Personal [4]. Philosopher [14, 15, 16, 17]. Photon [18]. Physics [12, 14, 15]. Physik [18]. Planck [2]. Pp [17, 16, 14, 15]. Preface [3]. production [13]. Proposal [18]. Published [4]. Publishing [16]. Quanta [2, 6]. Quantification [8]. Radiation [1]. Recollections [4]. Recorded [4]. Reevaluation [12]. Reference [11]. Reminiscences [4]. Review [17, 14, 15, 16]. Rutherford [4]. Sage [10]. Samuel [4]. School [12]. scientific [5]. Series [14, 15]. Table [11]. Theory [10]. Thought [14, 15, 5]. Time [4]. transformation [13]. Translation [18, 5, 3]. Verwandlung [13]. viewpoint [13]. vii [17]. Vol [17]. Volume [14, 15]. Volumes [14, 15]. Wave [9]. York [17, 14, 15, 16].

References

Beardsley:1963:RC

[1] Robert E. Beardsley. Radiation control. The Natural Philosopher, 1:23–??, 1963. CODEN NPHLAC. ISSN 0547-9592.

Klein:1963:PEQ

[2] Martin J. Klein. Planck, entropy, and quanta, 1901–1906. The Natural Philosopher, 1:83–108, 1963. CODEN NPHLAC. ISSN 0547-9592.

Anonymous:1963:NTP

[3] Anonymous. A translation of the preface to Lavoisier’s General Introduction to Chemistry (1787). The Natural Philosopher, 1:109–126, 1963. CODEN NPHLAC. ISSN 0547-9592.

Kay:1963:RRB

[4] William Alexander Kay. Recollections of Rutherford. Being the personal reminiscences of Lord Rutherford’s laboratory assistant. Here published for the first time. Recorded and annotated by Samuel Devons. The Natural Philosopher, 1:127–155, 1963. CODEN NPHLAC. ISSN 0547-9592.

Gershenson:1963:FCA

[5] Daniel E. Gershenson and Daniel A. Greenberg. The first chapter of Aristotle’s foundation of scientific thought (Metaphysica, Liber A). A new translation. The Natural Philosopher, 2:5–55, 1963. CODEN NPHLAC. ISSN 0547-9592.

Klein:1963:EFP

[6] Martin J. Klein. Einstein’s first paper on quanta. The Natural Philosopher, 2:57–86, 1963. CODEN NPHLAC. ISSN 0547-9592.

Dibner:1963:GVM

[7] Bern Dibner. The great van Marum electrical machine. The Natural Philosopher, 2:87–103, 1963. CODEN NPHLAC. ISSN 0547-9592.

King:1963:QCE

[8] W. James King. The quantification of the concepts of electric charge and electric current. Part I. The Natural Philosopher, 2:107–127, 1963. CODEN NPHLAC. ISSN 0547-9592.

Klein:1964:EWP

[9] Martin J. Klein. Einstein and the wave–particle duality. The Natural Philosopher, 3:3–49, 1964. CODEN NPHLAC. ISSN 0547-9592.

Aronson:1964:GTG

[10] Samuel Aronson. The gravitational theory of Georges-Louis Le Sage. The Natural Philosopher, 3:51–74, 1964. CODEN NPHLAC. ISSN 0547-9592.

Scott:1964:BRT

[11] William T. Scott. A bibliographic reference table for Faraday’s papers on electricity. The Natural Philosopher, 3:77–95, 1964. CODEN NPHLAC. ISSN 0547-9592.

Gershenson:1964:PES

[12] Daniel E. Gershenson and Daniel A. Greenberg. The “physics” of the Eleatic School: A reevaluation. The Natural Philosopher, 3:99–111, 1964. CODEN NPHLAC. ISSN 0547-9592.

Einstein:1905:EVL

[13] Albert Einstein. Über einen die Erzeugung und Verwandlung des Lichtes betreffenden heuristischen Gesichtspunkt. (German) [On the production and transformation of light from a heuristic viewpoint]. Annalen der Physik (1900) (series 4), 322(6):132–148, 1905. ISSN 0003-3804. URL http://

www.gsjournal.net/Science-Journals/Essays/View/2490;   http:/
/www.gsjournal.net/Science-Journals/Essays/View/2491;   http:
//www.zbp.univie.ac.at/einstein/einstein1.pdf. English translation

in [18].

Marton:1964:BRNa

[14] L. Marton. Book review: The Natural Philosopher, Volume 1. A Series of Volumes Devoted to the History of Physics and to the Influence of Physics on Human Thought and Affairs through the Ages, 155 pp., Blaisdell, New York, 1963. Physics Today, 17(6):56–57, June 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://scitation.aip.org/content/aip/magazine/
physicstoday/article/17/6/10.1063/1.3051656.

Marton:1964:BRNb

[15] L. Marton. Book review: The Natural Philosopher, Volume 2. A Series of Volumes Devoted to the History of Physics and to the Influence of Physics on Human Thought and Affairs through the Ages, 127 pp., Blaisdell, New York, 1963. Physics Today, 17(9):70–71, September 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://scitation.aip.org/docserver/fulltext/aip/magazine/
physicstoday/17/9/1.3051865.pdf.

Schmitt:1964:BRB

[16] Charles B. Schmitt. Book review: The Natural Philosopher 2, edited by Daniel E. Gershenson and Daniel A. Greenberg. New York: Blaisdell Publishing Company, 1963. Pp. 127. International Philosophical Quarterly, 4(3):493–494, December 1964. ISSN 0019-0365 (print), 2153-8077 (electronic).

Anderson:1965:RBN

[17] David L. Anderson. Review: The Natural Philosopher, Vol. III, edited by Daniel E. Gershenson and Daniel A. Greenberg. New York: Blaisdell, 1964, Pp. vii + 111. Technology and Culture, 6(3):536–537, Summer 1965. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

http://www.jstor.org/stable/3101839.

Arons:1965:EPP

[18] A. B. Arons and M. B. Peppard. Einstein’s proposal of the photon concept — a translation of the Annalen der Physik paper of 1905. American Journal of Physics, 33(5):367–374, May 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://scitation.aip.org/

content/aapt/journal/ajp/33/5/10.1119/1.1971542. English transla-

tion, with commentary in abstract, of Einstein’s 1905 paper [13] for which he won the 1921 Nobel Prize in Physics.